Question Statement and details : 500 mL of 56V H2O2 is kept in an open container due to which some H2O2 is decomposed and evolves 8gm O2 simultaneously, during the process some H2O also vapourizes.Due to all these changes final volume is reduced by 20%. Find final volume strength of H2O2 (aq). It's a multiple choice question. Options are as follows (a.) 56V (b) 44.8 V (c.) 11.2V (d.) 33.6V. 
 

Grade:11

1 Answers

Arun
4 years ago
 
Vol strength= 56V
Vol of O2 produced by 500ml solution = 56*500 = 28 Lt.
Moles of O2 produced = 28/22.4 = 1.25 mol
1mol of O2 produced by 2mol H2O2 
=> 1.25 mol of O2 produced by 2.5 mol of H2O2
So, mass of H2O2 = 2.5*34= 85g
Density= 170g/Lt
O2 escaped= 8g or 0.25 mol
Hence 0.5 mol of H2O2 decomposes to produce 0.5 mol of H2O n 0.25 mol of O2
Loss = 20% of 85g  = mass of O2 escaped+ mass of H2O vapourised
=> mass of H2O vapourised = 17 - 8 = 9g
So final solution contains only H2O2
Mass of final solution = 85 - 17 = 68g
Final vol = 68/170 = 400ml
68g H2O2 gives 32g O2
=> 400ml of solution gives 1mol of O2 or 22400ml O2
=> 1ml solution gives = 22400/400= 56ml of O2
Hence vol strength = 56V